1. Calculate the work done in kilojoules on a 1,146 kg elevator car by its cable to lift it 42 m at constant speed, assuming friction averages 135.9 N.

2. What is the cost of operating a 4.4 W electric clock for a year if the cost of electricity is $0.08 per kWh ?

3. Find the final speed for a skier in a downhill ski race who skies 80 m along a 24.7º slope neglecting friction if he or she starts from rest.
Answer #1

Here ,

for the elevator moving at constant speeed ,

tension in the cable , T = m * g + friction

T = 1146 * 9.8 + 135.9

work done = T * height

work done = (1146 * 9.8 + 135.9) * 42

work done = 477401.4 J

work done = 477.4 kJ

the work dobe by the cable is 477.4 kJ